OpenStudy (anonymous):

Can anyone just tell me what this means in interval notation? {x element R : 6 x+sqrt(33)<=0 or 6 x>=sqrt(33)} {x element R : 2 x>=1} {x element R : 32 x>=17}

OpenStudy (ybarrap):

I'll do the 1st for you: $$ 6x+\sqrt{33}\le 0\\ 6x \le -\sqrt{33}\\ x\le \cfrac{-\sqrt{33}}{6} $$ OR $$ 6 x>=\sqrt{33}\\ x\ge\cfrac{\sqrt{33}}{6} $$ Which means in interval notation: $$ (-\infty,\cfrac{-\sqrt{33}}{6}]\cup[x\ge\cfrac{\sqrt{33}}{6},\infty) $$

OpenStudy (ybarrap):

Correction: $$ \left (-\infty,\cfrac{-\sqrt{33}}{6}\right ]\cup\left [\cfrac{\sqrt{33}}{6},\infty\right ) $$